# Optimal e-value testing for properly constrained hypotheses

### v1

@@ Appendix A Omitted proofs
By Lemma 2, it is enough to show that every e-variable is majorised by a maximal e-variable, which implies that the set of all maximal e-variables is a majorising class. Fix $E_{0}\in\mathcal{E}_{\mathcal{H}}$. If $E_{0}$ is already maximal, we are done. So, we focus on the case of an $E_{0}$ that is not maximal.
Consider the transfinite sequence, starting from $E_{0}$ and indexed in $\omega_{1}$ (the smallest uncountable ordinal), defined as follows. For any countable successor ordinal $\beta\geq 1$, let $E_{\beta}=E_{\beta-1}$ if $E_{\beta-1}$ is maximal, otherwise pick any $E\in\mathcal{E}_{\mathcal{H}}$ such that $E\succ E_{\beta-1}$ and let $E_{\beta}=E$. For any limiting ordinal $\gamma\in\omega_{1}$, let $(\beta_{n})_{n\geq 1}\subseteq\omega_{1}$ be an increasing sequence of ordinals converging to $\gamma$ (with respect to the order topology). For $x\in\mathcal{X}$, let $A(x)=1/\sup_{P\in\mathcal{H}}P(\{x\})\in[1,+\infty)$. Note that for any $E\in\mathcal{E}_{\mathcal{H}}$ and any $x\in\mathcal{X}$, $E(x)=A(x)E(x)\sup_{P\in\mathcal{H}}P(\{x\})\leq A(x)\sup_{P\in\mathcal{H}}\langle P,E\rangle\leq A(x)$. So, for each $x\in\mathcal{X}$, the sequence $(E_{\beta_{n}}(x))_{n\geq 1}$ is non-decreasing and bounded in $[0,A(x)]$, and hence admits a finite limit. Therefore, we can define the point-wise limit $E_{\gamma}=\lim_{n\to\infty}E_{\beta_{n}}$. Moreover, by Beppo Levi theorem, $E_{\gamma}$ is Borel and $\langle P,E_{\gamma}\rangle\leq 1$, for any $P\in\mathcal{H}$. So, $E_{\gamma}\in\mathcal{E}_{\mathcal{H}}$.
Now, consider the set $B\subseteq\omega_{1}$ of the indices where the transfinite sequence has jumps (namely, $E_{\beta}\succ E_{\beta-1}$ if, and only if, $\beta\in B$). Since $E_{0}$ is not maximal, $B$ is non-empty. For each $\beta\in B$ we can select a pair $(x_{\beta},q_{\beta})\in\mathcal{X}\times\mathbb{Q}$ such that $E_{\beta-1}(x_{\beta})<q_{\beta}<E_{\beta}(x_{\beta})$. This defines an injection $B\to\mathcal{X}\times\mathbb{Q}$, whose existence implies that $B$ is countable, because $\mathcal{X}\times\mathbb{Q}$ is countable. Thus, $\beta^{\star}=\sup B\in\omega_{1}$, as it is the supremum of a countable subset of $\omega_{1}$. By construction, $E_{\beta^{\star}}\succeq E_{0}$ and $E_{\beta^{\star}}$ is maximal, since it must be that $E_{\beta^{\star}+1}=E_{\beta^{\star}}$, as otherwise $\beta^{\star}+1$ would be in $B$. ∎
Lemma 5.
Every properly constrained hypothesis admits a minimal constraint.
@@
Lemma 6.
Proof.
@@
Proposition 5.
[…]
Then, the optimal e-class for $\mathcal{H}$ exists and coincides with
Proof.
Let $\Phi:\mathcal{X}\to\mathbb{R}^{m^{\prime}+m^{\prime\prime}}$ be the continuous function whose first $m^{\prime}$ components are $\Phi^{\prime}$ and the last $m^{\prime\prime}$ are $\Phi^{\prime\prime}$. Since $\mathcal{H}$ is loosely properly constrained, $0\in\operatorname*{rel\,int}(\operatorname*{conv}\Phi)$. Moreover, note that $\mathcal{P}_{\Phi}=\mathcal{P}_{\Phi^{\prime}}\cap\mathcal{P}_{\Phi^{\prime\prime}}$, so =$\mathcal{H}^{\prime}=\{P\in\mathcal{P}_{\Phi^{\prime}}\cap\mathcal{P}_{\Phi^{\prime\prime}}\,:\,\langle P,\Phi\rangle=0\}$ is a properly constrained hypothesis, whereof $\Phi$ is a constraint. By construction $\mathcal{H}^{\prime}\subseteq\mathcal{H}$, so $\mathcal{E}_{\mathcal{H}^{\prime}}\supseteq\mathcal{E}_{\mathcal{H}}$. Moreover, $\mathcal{E}_{\mathcal{H}^{\prime}}^{\vee}\supseteq\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}$. In particular, since every $E\in\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}$ is a maximal e-variable for $\mathcal{H}^{\prime}$, it must also be a maximal e-variable for $\mathcal{H}$. Hence, by Lemma 2 it is enough to show that $\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}$ is a majorising e-class. This can be done reproducing step by step the proofs of Proposition 4 and Theorem 1. Indeed, the preliminary results that were used for those proofs are essentially based on the compactness of $\Lambda_{\Phi|S}$ (for suitable choices of $S\subseteq\mathcal{X}$). This property holds also for $\tilde{\Lambda}_{\Phi^{\prime},\Phi^{\prime\prime}|S}$ (with obvious meaning of notation), as $\tilde{\Lambda}_{\Phi^{\prime},\Phi^{\prime\prime}|S}$ is a closed subset of $\Lambda_{\Phi|S}$ (for the same argument as in Lemma 16). ∎
@@ B.1 Some linear algebraic results
Let $\{v_{1},\dots,v_{r}\}\subseteq\mathbb{R}^{m}$ be a family of vectors that spans $\mathbb{R}^{m}$. Let $\{(A_{i},B_{i})\}_{i=1}^{r}$ be a family of non-empty bounded intervals. Then, the set
$\Lambda=\{\lambda\in\mathbb{R}^{m}\;:\;\lambda\cdot v_{i}\in[A_{i},B_{i}]\,,\;\forall i=1\dots r\}$
is bounded.
@@
Since $\{v_{1},\dots,v_{r}\}$ spans $\mathbb{R}^{m}$, each element $e_{i}$ of the standard basis of $\mathbb{R}^{m}$ can be written as a linear combination $e_{i}=\sum_{j=1}^{r}\alpha_{ij}v_{j}$. Then, if $\lambda\cdot v_{j}$ is in a bounded interval for each $j$, it follows that $\lambda\cdot e_{i}$ is also in a bounded interval. Hence, all the components of $\lambda$ must be bounded if $\lambda\in\Lambda$, and so $\Lambda$ is bounded. ∎
@@ B.2 Some results from convex analysis
First, we state some standard definitions for elementary notions in convex analysis. We refer to Hiriart-Urrut and Lemaréchal (2004) for a thorough introduction to the subject. Given a set $\mathcal{Y}\subseteq\mathbb{R}^{n}$, its convex hull $\mathcal{C}=\operatorname*{conv}\mathcal{Y}$ is defined as the smallest convex set containing $\mathcal{Y}$, or equivalently, as the intersection of all the convex sets that contain $\mathcal{Y}$. The affine hull $\mathcal{A}=\operatorname*{aff}\mathcal{Y}$ of $\mathcal{Y}$ is the smallest affine space in $\mathbb{R}^{n}$ that contains $\mathcal{Y}$, while the linear hull (or span) $\mathcal{L}=\operatorname*{lin}\mathcal{Y}$ of $\mathcal{Y}$ is the smallest vector subspace of $\mathbb{R}^{n}$ containing $\mathcal{Y}$. Clearly,
$\mathcal{Y}\subseteq\mathcal{C}\subseteq\mathcal{A}\subseteq\mathcal{L}\,.$
Let $C\subseteq\mathbb{R}^{n}$ be a convex set and $\mathcal{A}$ its affine hull. The relative interior of $C$ (denoted as $\operatorname*{rel\,int}C$) is the interior of $C$ with respect to the induced topology on $\mathcal{A}$. More explicitly, $x\in C$ is in $\operatorname*{rel\,int}C$ if, and only if, there is an open set $U$ in $\mathbb{R}^{n}$ containing $x$ and such that $U\cap\mathcal{A}\subseteq C$. The relative boundary of $C$ is the set $C\setminus\operatorname*{rel\,int}C$.
@@
Let $m^{\prime}$ be the dimension of $\operatorname*{Span}\Phi$. We can find $m^{\prime}$ components of $\Phi$ that are linearly independent. Without loss of generality, we assume that these are the first $m^{\prime}$ components, and let $\Phi^{\prime}=(\Phi_{1},\dots,\Phi_{m^{\prime}})$, which is a function $\mathcal{X}\to\mathbb{R}^{m^{\prime}}$ with $\operatorname*{Span}\Phi^{\prime}$ of dimension $m^{\prime}$. By Lemma 20, there is a family $S=\{x_{1},\dots,x_{m^{\prime}}\}$ of $m^{\prime}$ distinct points in $\mathcal{X}$ such that the $m^{\prime}\times m^{\prime}$ Gram matrix with components $\Phi_{i}(x_{j})$ (for $i$ and $j$ ranging from $1$ to $m^{\prime}$) has full rank. In particular, $\operatorname*{lin}\Phi^{\prime}(S)=\mathbb{R}^{m^{\prime}}$. Since $\mathbb{R}^{m^{\prime}}\supseteq\operatorname*{lin}\Phi^{\prime}(\mathcal{X})\supseteq\operatorname*{lin}\Phi^{\prime}(S)$, we conclude that $\dim(\operatorname*{lin}\Phi^{\prime}(\mathcal{X}))=m^{\prime}$. So, $\dim(\operatorname*{lin}\Phi(\mathcal{X}))\geq m^{\prime}$. Now, Lemma 20 also implies that no subset of $\mathcal{X}$ with strictly more than $m^{\prime}$ elements can yield a fully ranked Gram matrix. Therefore, there cannot be more than $m^{\prime}$ independent vectors in $\operatorname*{lin}\Phi(\mathcal{X})$, and so $\dim(\operatorname*{lin}\Phi(\mathcal{X}))=m^{\prime}$. ∎
@@
Let $\mathbb{R}^{d}_{\geq 0}$ denote the subset of $\mathbb{R}^{d}$ of vectors whose components are all non-negative. Its relative interior is >$\mathbb{R}^{d}_{>0}$, the set of vectors with only strictly positive components. Let $\mathbf{1}\in\mathbb{R}^{d}$ denote the vector with all components equal to one and $V=\{v\in\mathbb{R}^{d}\,:\,v\cdot\mathbf{1}=1\}$, whose relative interior is $V$ itself. Then $\mathcal{P}_{\mathcal{X}}=\mathbb{R}^{d}_{\geq 0}\cap V$, which is the non-empty intersection of two convex sets. Since $\operatorname*{rel\,int}V\cap\operatorname*{rel\,int}\mathbb{R}^{d}_{\geq 0}\neq\varnothing$, from Proposition 2.1.10 in Chapter A of Hiriart-Urrut and Lemaréchal, 2004 we get
@@
Appendix C An example of non-existence of the optimal e-class
Let $\mathcal{X}=[0,1]$. Denote as $V$ any non-Borel set in $[0,1]$, which contains 00 and does not contain $1$. For $x\in[0,1]$, let $P_{x}$ denote $\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{x}$. Define $\mathcal{H}$ as follows
=$\mathcal{H}=\big{\{}P_{x}\;:\;x\in[0,1]\big{\}}\cup\mathcal{P}_{V}\,.$
Recall that $\mathcal{P}_{V}$ denote the set of all probability measures in $[0,1]$ whose support lies in $V$. So, even if $V$ is non-measurable, $\mathcal{P}_{V}$ is well defined. Clearly, $\mathcal{H}$ is a hypothesis on $[0,1]$, and
$\sup_{P\in\mathcal{H}}P(\{x\})=\begin{cases}1&\text{if $x\in V$;}\\ 1/2&\text{otherwise.}\end{cases}$
Now, define the function $A:[0,1]\to\mathbb{R}$ as $A(x)=1/\sup_{P\in\mathcal{H}}P(\{x\})$. Note that $A$ is not Borel measurable, so $A\notin\mathcal{E}_{\mathcal{H}}$. For any $E\in\mathcal{E}_{\mathcal{H}}$ and $x\in[0,1]$, we have
$E(x)\leq A(x)E(x)\sup_{P\in\mathcal{H}}P(\{x\})\leq A(x)\sup_{P\in\mathcal{H}}\langle P,E\rangle\leq A(x)\,.$
We will show that every e-variable $E$ such that $E(1)=2$ cannot be maximal. Since every majoriser $\hat{E}$ of such $E$ has to still satisfy $\hat{E}(1)=2$ (as $22=E(1)\leq\hat{E}(1)\leq A(1)=$ 2), we will then conclude that $E$ does not have any maximal majoriser and so that the class of all maximal e-variables is not a majorising class. Lemma 2 will yield that the optimal e-class does not exist.
So, fix an e-variable $E$ such that $E(1)=2$ (such an e-variable exists, for instance one can pick the function that is $2$ on $1$ and 00 everywhere else). Since $\langle P_{1},E\rangle\leq 1$, it must be that $E(0)=0$. The restriction of $A$ to $(0,1]$ is non-Borel, so there must be $x_{0}\in(0,1]$ such that $E(x_{0})<A(x_{0})$. Define $E^{\prime}:[0,1]\to[0,+\infty)$ as $E^{\prime}(x)=E(x)$, for $x\neq x_{0}$, and $E^{\prime}(x_{0})=A(x_{0})$. Since $E^{\prime}\preceq A$, we have that $E^{\prime}(x)\leq 1$ for any $x\in V$. In particular, for any $P\in\mathcal{P}_{V}$ we have $\langle P,E^{\prime}\rangle\leq 1$. Moreover, $E^{\prime}(0)=E(0)=0$ and $E^{\prime}(x)\preceq 2$ for all $x\in[0,1]$. Thus, $\langle P_{x},E^{\prime}\rangle=\frac{1}{2}E^{\prime}(x)\leq 1$, for any $x\in[0,1]$. We deduce that $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$. Since $E^{\prime}\succ E$, $E$ is not maximal, and so we conclude.
Appendix D Proof of Proposition 1

### v3

@@ Acknowledgements
Since the initial appearance of this manuscript, parts of it have been simplified thanks to valuable suggestions. In particular, I would like to thank Martin Larsson for proposing a much simpler example demonstrating the non-existence of the optimal e-class, which, unlike the example in the original version, avoids the use of non-measurable sets.
@@ Appendix A Omitted proofs
This proof makes use of elementary properties of ordinals. For an accessible reference, see Aliprantis and Border (2006).
By Lemma 1, it is enough to show that every e-variable is majorised by a maximal e-variable, which implies that the set of all maximal e-variables is a majorising class. Fix $E_{1}\in\mathcal{E}_{\mathcal{H}}$. If $E_{1}$ is already maximal, we are done. So, we focus on the case of an $E_{1}$ that is not maximal.
Consider the transfinite sequence (namely, net), starting from $E_{1}$ and indexed in $\Omega_{0}=[1,\omega_{1})$ (the set of the countable ordinals), defined as follows. For any successor ordinal $\beta>1$ in $\Omega_{0}$, let $E_{\beta}=E_{\beta-1}$ if $E_{\beta-1}$ is maximal, otherwise pick any $E\in\mathcal{E}_{\mathcal{H}}$ such that $E\succ E_{\beta-1}$ and let $E_{\beta}=E$. For any limiting ordinal $\gamma\in\Omega_{0}$, let $(\beta_{n})_{n\geq 1}\subseteq\Omega_{0}$ be an increasing sequence converging to $\gamma$ (under the order topology). For $x\in\mathcal{X}$, let $A(x)=1/\sup_{P\in\mathcal{H}}P(\{x\})\in[1,+\infty)$. Note that for any $E\in\mathcal{E}_{\mathcal{H}}$ and any $x\in\mathcal{X}$, $E(x)=A(x)E(x)\sup_{P\in\mathcal{H}}P(\{x\})\leq A(x)\sup_{P\in\mathcal{H}}\langle P,E\rangle\leq A(x)$. So, for each $x\in\mathcal{X}$, the sequence $(E_{\beta_{n}}(x))_{n\geq 1}$ is non-decreasing and bounded in $[0,A(x)]$, and hence admits a finite limit. Therefore, we can define the point-wise limit $E_{\gamma}=\lim_{n\to\infty}E_{\beta_{n}}$. Moreover, by Beppo Levi theorem, $E_{\gamma}$ is Borel and $\langle P,E_{\gamma}\rangle\leq 1$, for any $P\in\mathcal{H}$. So, $E_{\gamma}\in\mathcal{E}_{\mathcal{H}}$.
Now, consider the set $B\subseteq\Omega_{0}$ of the indices where the transfinite sequence has jumps (namely, $E_{\beta}\succ E_{\beta-1}$ if, and only if, $\beta\in B$). Since $E_{1}$ is not maximal, $B$ is non-empty. For each $\beta\in B$ we can select a pair $(x_{\beta},q_{\beta})\in\mathcal{X}\times\mathbb{Q}$, such that $E_{\beta-1}(x_{\beta})<q_{\beta}<E_{\beta}(x_{\beta})$. This defines an injection $B\to\mathcal{X}\times\mathbb{Q}$, whose existence implies that $B$ is countable, because $\mathcal{X}\times\mathbb{Q}$ is countable. As every countable subset of $\Omega_{0}$ has a least upper bound in $\Omega_{0}$ (Theorem 1.14 in Aliprantis and Border, 2006), we have $\beta^{\star}=\sup B\in\Omega_{0}$. By construction, $E_{\beta^{\star}}\succeq E_{1}$ and $E_{\beta^{\star}}$ is maximal, since it must be that $E_{\beta^{\star}+1}=E_{\beta^{\star}}$, as otherwise $\beta^{\star}+1$ would be in $B$. ∎
Proposition 2.
Let $\mathcal{X}=[0,1]$. Consider the hypothesis
=$\mathcal{H}=\big{\{}P\in\mathcal{P}_{[0,1]}\,:\,P(\{0\})\geq 1/2\big{\}}\cup\{U_{[0,1]}\}\,,$
with $U_{[0,1]}$ the uniform distribution on $[0,1]$. $\mathcal{H}$ does not admit an optimal e-class.
@@
First, let us notice that all e-variables must be everywhere upper bounded by $2$. Indeed, for any $x\in[0,1]$ we have that $P_{x}=\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{x}$ is in $\mathcal{H}$. In particular, for any e-variable $E$ we have $\langle P_{x},E\rangle=\frac{1}{2}(E(x)+E(0))\geq\frac{1}{2}E(x)$, and so $E(x)\leq 2$.
Let $\mathcal{E}_{1}$ denote the set of all e-variables such that $E(1)=2$. Clearly, this is non-empty, since the function that is 00 everywhere and $2$ on $1$ belongs to $\mathcal{E}_{1}$. We now show that no variable in $\mathcal{E}_{1}$ can be maximal. As a consequence, since all e-variables in $\mathcal{E}_{1}$ can only be majorised by e-variables in $\mathcal{E}_{1}$, we conclude that there is no maximal e-variable that majorises the e-variables in $\mathcal{E}_{1}$, and so the set of all maximal e-variables is not a majorising e-class. By Lemma 1 we conclude that the optimal e-class does not exist.
Now, let $E\in\mathcal{E}_{1}$. First note that $E(0)=0$, as with $P_{1}=\frac{1}{2}\delta_{0}+\frac{1}{2}\delta 1$ we have that $1\geq\frac{1}{2}(E(0)+E(1))=1+\frac{1}{2}E(0)$. Moreover, since $\langle U_{[0,1]},E\rangle\leq 1$, $E$ cannot be identically equal to $2$ and there must be at least one $x^{\prime}\in(0,1]$ such that $E(x^{\prime})<2$. Define $E^{\prime}$ as
$E^{\prime}(x)=\begin{cases}E(x)&\text{if $x\neq x^{\prime}$;}\\ 2&\text{if $x=x^{\prime}$.}\end{cases}$
Clearly, $E^{\prime}\succ E$. So, it is enough to show that $E^{\prime}$ is an e-variable to establish that $E$ cannot be maximal.
Fix any $P$ such that $P(\{0\})\geq 1/2$. We can write $P=P(\{0\})\delta_{0}+(1-P(\{0\}))P^{\prime}$, where $P^{\prime}\in\mathcal{P}_{[0,1]}$ is such that $P^{\prime}(\{0\})=0$. We have
$\langle P,E^{\prime}\rangle=P(\{0\})E^{\prime}(0)+(1-P(\{0\}))\langle P^{\prime},E^{\prime}\rangle\,.$
Since $x^{\prime}>0$, $E^{\prime}(x)=E(x)=0$, so $\langle P,E^{\prime}\rangle\leq\frac{1}{2}\langle P^{\prime},E^{\prime}\rangle\leq 1$, since $E^{\prime}\preceq 2$. We are only left with checking that $\langle U_{[0,1]},E^{\prime}\rangle\leq 1$. But since $E$ and $E^{\prime}$ differs only on $x^{\prime}$, it is clear that $\langle U_{[0,1]},E^{\prime}\rangle=\langle U_{[0,1]},E\rangle$, and so we conclude. ∎
@@
Lemma 6.
Proof.
Lemma 7.
Every properly constrained hypothesis admits a minimal constraint.
Proof.
@@ B.1 Some linear algebraic results
Let $\{v_{1},\dots,v_{r}\}\subseteq\mathbb{R}^{m}$ be a family of vectors that spans $\mathbb{R}^{m}$. Let $\{(A_{i},B_{i})\}_{i=1}^{r}$ be a family of non-empty bounded intervals. Then, the set $\Lambda=\{\lambda\in\mathbb{R}^{m}\;:\;\lambda\cdot v_{i}\in[A_{i},B_{i}]\,,\;\forall i=1\dots r\}$ is bounded.
@@
Since $\{v_{1},\dots,v_{r}\}$ spans $\mathbb{R}^{m}$, each element $e_{i}$ of the standard basis of $\mathbb{R}^{m}$ can be written as a linear combination $e_{i}=\sum_{j=1}^{r}\alpha_{ij}v_{j}$. Then, if $\lambda\cdot v_{j}$ is in a bounded interval for each $j$, it follows that $\lambda\cdot e_{i}$ is also in a bounded interval. Hence, all the components of $\lambda$ must be bounded if $\lambda\in\Lambda$. ∎
@@ B.2 Some results from convex analysis
First, we state some standard definitions for elementary notions in convex analysis. We refer to Hiriart-Urrut and Lemaréchal (2004) for a thorough introduction to the subject. Given a set $\mathcal{Y}\subseteq\mathbb{R}^{n}$, its convex hull $\mathcal{C}=\operatorname*{conv}\mathcal{Y}$ is defined as the smallest convex set containing $\mathcal{Y}$, or equivalently, as the intersection of all the convex sets that contain $\mathcal{Y}$. The affine hull $\mathcal{A}=\operatorname*{aff}\mathcal{Y}$ of $\mathcal{Y}$ is the smallest affine space in $\mathbb{R}^{n}$ that contains $\mathcal{Y}$, while the linear hull (or span) $\mathcal{L}=\operatorname*{lin}\mathcal{Y}$ of $\mathcal{Y}$ is the smallest vector subspace of $\mathbb{R}^{n}$ containing $\mathcal{Y}$. Clearly, $\mathcal{Y}\subseteq\mathcal{C}\subseteq\mathcal{A}\subseteq\mathcal{L}$. Let $C\subseteq\mathbb{R}^{n}$ be a convex set and $\mathcal{A}$ its affine hull. The relative interior of $C$ (denoted as $\operatorname*{rel\,int}C$) is the interior of $C$ with respect to the induced topology on $\mathcal{A}$. More explicitly, $x\in C$ is in $\operatorname*{rel\,int}C$ if, and only if, there is an open set $U$ in $\mathbb{R}^{n}$ containing $x$ and such that $U\cap\mathcal{A}\subseteq C$. The relative boundary of $C$ is the set $C\setminus\operatorname*{rel\,int}C$.
@@
Let $m^{\prime}$ be the dimension of $\operatorname*{Span}\Phi$. We can find $m^{\prime}$ components of $\Phi$ that are linearly independent. Without loss of generality, we assume that these are the first $m^{\prime}$ components, and let $\Phi^{\prime}=(\Phi_{1},\dots,\Phi_{m^{\prime}})$, which is a function $\mathcal{X}\to\mathbb{R}^{m^{\prime}}$ with $\operatorname*{Span}\Phi^{\prime}$ of dimension $m^{\prime}$. By Lemma 19, there is a family $S=\{x_{1},\dots,x_{m^{\prime}}\}$ of $m^{\prime}$ distinct points in $\mathcal{X}$ such that the $m^{\prime}\times m^{\prime}$ Gram matrix with components $\Phi_{i}(x_{j})$ (for $i$ and $j$ ranging from $1$ to $m^{\prime}$) has full rank. In particular, $\operatorname*{lin}\Phi^{\prime}(S)=\mathbb{R}^{m^{\prime}}$. Since $\mathbb{R}^{m^{\prime}}\supseteq\operatorname*{lin}\Phi^{\prime}(\mathcal{X})\supseteq\operatorname*{lin}\Phi^{\prime}(S)$, we conclude that $\dim(\operatorname*{lin}\Phi^{\prime}(\mathcal{X}))=m^{\prime}$. So, $\dim(\operatorname*{lin}\Phi(\mathcal{X}))\geq m^{\prime}$. Lemma 19 also implies that no subset of $\mathcal{X}$ with strictly more than $m^{\prime}$ elements yields a fully ranked Gram matrix. So, there are at most $m^{\prime}$ independent vectors in $\operatorname*{lin}\Phi(\mathcal{X})$, and so $\dim(\operatorname*{lin}\Phi(\mathcal{X}))=m^{\prime}$. ∎
@@
Let $\mathbb{R}^{d}_{\geq 0}$ denote the subset of $\mathbb{R}^{d}$ of vectors whose components are all non-negative. Its relative interior is >$\mathbb{R}^{d}_{>0}$, the set of vectors with only strictly positive components. Let $\mathbf{1}\in\mathbb{R}^{d}$ denote the vector with all components equal to one and $V=\{v\in\mathbb{R}^{d}\,:\,v\cdot\mathbf{1}=1\}$, whose relative interior is $V$ itself. Then $\mathcal{P}_{\mathcal{X}}=\mathbb{R}^{d}_{\geq 0}\cap V$, which is the non-empty intersection of two convex sets. Since $\operatorname*{rel\,int}V\cap\operatorname*{rel\,int}\mathbb{R}^{d}_{\geq 0}\neq\varnothing$, from Proposition 2.1.10 in Chapter A of Hiriart-Urrut and Lemaréchal (2004), we get
@@
Appendix C Proof of Proposition 6
Here we provide a proof of Proposition 6. First, we discuss the case of $\mathcal{X}$ with finite cardinality. Much of the proof follows that of Lemma 2, but for the final step a different path needs to be taken as it is not anymore true that $\langle P,E\rangle=1$ for any $P\in\mathcal{H}$ if $E$ is maximal.
We start by an easy lemma.